A piston cylinder device with a free floating piston contains 0.3 kg of R-134a at 400 kPa and 0 C. The volume of the cylinder is what?
Answer #1

MOlecular mass of R-134a = 102.3 g / mol

number of moles = W/M = 300/102.3 = 2.93

Using the Universal gas equation we have

PV = nRT

(400000 / 1.01235*10^5) * V = 2.93*0.0821*273

V = 16.63 L

volume of cylinder is 16.63 L